# StreamVoice: Streamable Context-Aware Language Modeling for Real-time Zero-Shot Voice Conversion

### v1

@@ Abstract
Recent language model (LM) advancements have showcased impressive zero-shot voice conversion (VC) performance. However, existing LM-based VC models usually apply offline conversion from source semantics to acoustic features, demanding the complete source speech, and limiting their deployment to real-time applications. In this paper, we introduce StreamVoice, a novel streaming LM-based model for zero-shot VC, facilitating real-time conversion given arbitrary speaker prompts and source speech. Specifically, to enable streaming capability, StreamVoice employs a fully causal context-aware LM with a temporal-independent acoustic predictor, while alternately processing semantic and acoustic features at each time step of autoregression which eliminates the dependence on complete source speech. To address the potential performance degradation from the incomplete context in streaming processing, we enhance the context-awareness of the LM through two strategies: 1) teacher-guided context foresight, using a teacher model to summarize the present and future semantic context during training to guide the model’s forecasting for missing context; 2) semantic masking strategy, promoting acoustic prediction from preceding corrupted semantic and acoustic input, enhancing context-learning ability. Notably, StreamVoice is the first LM-based streaming zero-shot VC model without any future look-ahead. Experimental results demonstrate StreamVoice’s streaming conversion capability while maintaining zero-shot performance comparable to non-streaming VC systems.
Figure 1: The concept of the streaming zero-shot VC employing the widely used recognition-synthesis framework. StreamVoice is built on this popular paradigm.
@@ 1 Introduction
Voice conversion (VC) aims to transfer a speaker’s voice to that of another speaker without changing the linguistic content. Such technique has been deployed in many real-world applications, e.g., movie dubbing, privacy protection, and pronunciation correction etc. With the help of neural semantic features, such as the bottleneck feature (BNF) from an automatic speech recognition (ASR) system, converting source speech from arbitrary speakers in the wild has been successfully achieved Sun et al. (2016). Meanwhile, converting to an arbitrary target speaker with only one utterance of this speaker, which is the so-called zero-shot VC, also has been researched recently Qian et al. (2019); Wang et al. (2023c). However, most existing zero-shot VC models are designed for offline systems, which are insufficient to meet the recent growing demands of streaming capability in real-time VC applications, such as live broadcasting and real-time communication (RTC). In this study, we focus on the streaming zero-shot VC as illuminated in Fig. 1.
@@
Disentangling speech into different components, e.g., semantic content and speaker timbre, plays an important role in the zero-shot VC task Chou and Lee (2019); Wang et al. (2023d, 2021); Qian et al. (2019). Recently, benefiting from the powerful LM framework and the scaling up of training data, LM-based VC models Wang et al. (2023c); Yang et al. (2023); Zhu et al. (2023) with built-in in-context learning ability can sense the context relations between source and target speaker’s utterances to capture fine-grained speaker timbre, achieving impressive zero-shot VC performance. However, demanding the complete source speech utterance limits these LM-based VC models in real-time scenarios, and thus these models can only be used in offline applications. While several non-LM-based methods Yang et al. (2022); Wang et al. (2023a) have been proposed for streaming zero-shot VC, the performance fails to generalize well to unseen speakers with high speaker similarity and speech naturalness, mainly due to the limited model capacity to scale up training data, and also the performance degradation caused by the missing future information in streaming scenario.
@@
Inspired by the success of LM-based models in zero-shot VC, we aim to explore the feasibility of LMs for the streaming VC scenario. An intuitive way is to follow the popular recognize-synthesis framework as shown in Fig. 1, in which, speech is represented in semantic BNF and acoustic features respectively extracted by a streaming ASR and an audio codec. Then the LM-based VC model undertakes the transformation of semantic information into acoustic features with the target speaker’s timbre. However, the development of the LM-based model in streaming zero-shot VC is hampered by two primary challenges.
@@
  • Streamable architecture: streaming models typically produce immediate output upon receiving current input without reliance on future time steps. Current LM-based VC models perform the conversion only when get a full-utterance of source speech, which fails to meet the demands of streaming applications. The widely adopted multi-stage language modeling for multi-layer codec prediction introduces complexity to system design, thereby posing a potential risk of cumulative errors. Additionally, the dependency models of the streaming pipeline also impact the design and performance of the VC model.
@@
  • Performance gap: unlike non-streaming models, streaming models must process frame-wise or chunked input causally on the fly without future information, facing missing context and potential performance degradation. This missing hinders the streaming VC model from achieving high-quality conversion. In addition, as shown in Fig. 1, the VC model relies on the semantic feature BNF from ASR to achieve conversion, which makes semantic features very important. However, streaming ASR exhibits inferior performance compared to its non-streaming counterpart, leading to the extracted BNF carrying low-quality semantic information but more speaker information. In addition to the inherent unavailable future reception, this low-quality semantic input makes achieving high-quality conversion more difficult. The goal of zero-shot VC amplifies the challenges faced by our streaming VC model.
@@
In this work, we propose StreamVoice, a streaming LM-based model for high-quality zero-shot VC, in which teacher-guided semantic foresight and semantic masking are integrated to enhance the context awareness of the model for improving conversion quality. Specifically, StreamVoice creates a streamable architecture by integrating a single-stage language model that casually generates acoustic codecs with the collaboration of an acoustic predictor. Alternating input of semantic and acoustic features at each time step ensures seamless streaming behavior. To mitigate the performance gap caused by missing contextual information, two methods are introduced to enhance the context-awareness of the LM. 1) We incorporate a teacher-guided semantic foresight, where the VC model is taught by a teacher non-streaming ASR model to infer the present and future semantic information summarized by the teacher, which is then used to enhance the acoustic prediction. 2) To enhance the context learning from the input history, semantic masking is used to encourage acoustic prediction from the preceding acoustic and corrupted semantic input, which also implicitly creates an information bottleneck to reduce the source speaker’s information.
@@
Experiments demonstrate StreamVoice’s ability to convert speech in a streaming manner with high speaker similarity for both seen and unseen speakers while maintaining performance comparable to non-streaming VC systems. As the first LM-based zero-shot VC model without any future look-ahead, the total pipeline of StreamVoice only has 124 ms latency to perform the conversion, 2.4x faster than real-time on a single A100 GPU without engineering optimizations.
@@ 2.1 Zero-shot Voice Conversion
Zero-shot VC imposes stringent demands on speech decoupling and capturing speaker timbre. Many studies specifically design many disentanglement approaches, incorporating intricate structures Chou and Lee (2019), loss functions Wang et al. (2021), and training strategies Ebbers et al. (2021), to achieve speech decoupling. Rather than embedding explicit disentanglement designs in VC training, some approaches Gu et al. (2021) leverage a speaker verification (SV) model for speaker representation, while linguistic content is extracted using ASR or self-supervised learning (SSL) models Sun et al. (2016); Choi et al. (2021). To enhance speaker timbre capturing, some fine-grained speaker modeling methods also have been explored Yin et al. (2021); Wang et al. (2023d). Recent successes of language models in generative tasks have prompted the exploration of LM-based models in zero-shot VC, yielding impressive results. Using the pre-trained model to decouple speech, the LM-based VC model Wang et al. (2023c); Yang et al. (2023); Zhu et al. (2023) can capture fine-grained speaker timbre from the speaker prompt and then perform the conversion. However, current LM-based VC models are inapplicable to streaming scenarios, constraining their real-world utility. This paper addresses this gap by investigating the zero-shot capabilities of language models specifically tailored for streaming scenarios.
2.2 Streaming Voice Conversion
@@
Despite the high-quality conversion achieved by non-streaming VC models, their non-streamable structure and reliance on full-utterance input hamper them for real-time streaming applications. For streaming, causal processing and the structure of the streaming pipeline are crucial considerations. Streaming models are compelled to process frame-wise or chunked input on the fly, devoid of access to future information, leading to performance degradation compared to non-streaming counterparts. To address this, a common approach Hayashi et al. (2022); Kameoka et al. (2021); Ning et al. (2023) involves the integration of a teacher model to guide the training of the streaming model or the distillation of knowledge from a non-streaming model. Chen et al. Chen et al. (2023b) focus on selecting BNF with minimal semantic information loss through layer-wise analysis, while Chen et al. Chen et al. (2022) incorporate adversarial training to enhance the quality of semantic features. Beyond streaming VC, recently, there have also been some efforts towards streaming zero-shot VC. For instance, VQMIVC Wang et al. (2021), which is designed for the non-streaming application, is modified to be streamable by Yang et al. Yang et al. (2022). ALO-VC Wang et al. (2023a) constructs a streaming system using an SV model, a streaming PPG extractor, and a pitch extractor. However, current streaming zero-shot VC, designed for low-resource devices, has limited model capacity with poor generalization to unseen speakers, leading to inferior similarity and naturalness. Motivated by LM’s successes in zero-shot VC, we design a streamable LM in streaming scenarios. To tackle distinctive challenges in streaming VC, we introduce teacher-guided semantic foresight and semantic masking to enhance LM’s context awareness and improve conversion quality.
2.3 Language Model-based Speech Generation
@@
In recent years, advancements in language models (LMs) within natural language processing have showcased potent generation capabilities, influencing the development of LMs in speech generation. By employing codec Zeghidour et al. (2021) or other SSL models Chung et al. (2021), speech and audio can be efficiently tokenized into discrete units, facilitating low-bitrate audio representation and semantic information extraction. This progress allows speech generation to seamlessly utilize LM frameworks. Taking audio generation as a conditional language modeling task, AudioLM Borsos et al. (2023) and MusicLM Agostinelli et al. (2023) employ hierarchical language modeling for acoustic prediction from coarse to fine units. VALL-E Wang et al. (2023b) and SpearTTS Kharitonov et al. (2023) extend LMs for zero shot-TTS, which can clone a human’s voice with prompt tokens from a short recording. For zero-shot VC, LM-VC Wang et al. (2023c) employs task-oriented optimizations to this task. And some studies Zhu et al. (2023); Yang et al. (2023) leverage multitask objectives and datasets, achieving high-quality conversion. Despite this progress, existing LM-based VC models usually apply offline processing, demanding complete utterance from the source speech, which hinders their suitability for real-time streaming applications. In contrast to prior studies, we explore the zero-shot capability of the LM-based VC for streaming scenarios. With the enhancement of context awareness, the proposed LM-based VC model achieves results comparable to non-streaming LM-based VC.
@@ 3.1 Overview
As shown in Fig. 2, the development of StreamVoice follows the recognition-synthesis framework. In this framework, speech is first represented as semantic features $\mathbf{s}=\{s_{1},s_{2},...s_{T_{s}}\}$ and acoustic features $\mathbf{a}=\{a_{1},a_{2},...,a_{T_{a}}\}$ by a pre-trained streaming ASR model and a speech codec model respectively. Here, $T_{s}$ and $T_{a}$ denote the sequence length. Before inputting to StreamVoice, $\mathbf{s}$ and $\mathbf{a}$ are aligned to the same length $T$. StreamVoice incorporates a context-aware language model and an acoustic predictor to perform a single language modeling process. With the semantic and acoustic features $\{\mathbf{\tilde{s}},\mathbf{\tilde{a}}\}$ of speech from the target speaker as speaker prompt, the LM leverages the semantic information $\mathbf{a}_{1:t}$ of source speech to autoregressively predict the hidden output ${}^{c}\mathbf{h}$. In each autoregression time-step of the LM, the acoustic predictor transforms the hidden output ${}^{c}\mathbf{h}$ to the codec feature $\mathbf{\hat{a}}$ of the converted speech. Finally, the decoder of the codec model reconstructs the waveform from the predicted codec feature. In the following sections, we will introduce how to build a streamable LM for VC and how to ensure the high-quality conversation of this streaming VC.
Figure 3: The architecture for context-aware language model.
@@ 3.2 Streamable Architecture
To perform streaming voice conversion, a streamable architecture is necessary. In StreamVoice, the LM is carefully designed to perform a fully causal processing in the VC task, and the acoustic predictor is designed to achieve frame-wise prediction without dependency on temporal information.
@@ 3.2.1 Fully Casual Language Model
Figure 4: The architecture for acoustic predictor. Our system can support continuous or discrete codec projection.
@@ 3.2.2 Acoustic Predictor
As the preceding LM has essentially encoded content and speaker into its output ${}^{c}\mathbf{h}$, the acoustic predictor can be designed in temporal irrelevant to transform ${}^{c}\mathbf{h}$ into acoustic codec space, which means the predictor can be easily applied in the streaming scenario. Given that the speech can be represented in acoustic features by neural codec in either continuous or discrete forms, we investigate the incorporation of both features in StreamVoice, which are performed by continuous projection and discrete projection respectively.
Continuous Projection: Following Shen et al. Shen et al. (2023), the $D$-dimensional quantized latent vector $\mathbf{a}\in\mathcal{R}^{T\times D}$ encoded by the codec model is used as the continuous acoustic representation. The prediction of the continuous representation involves employing a stack of linear layers, as shown in Fig. 4. The continuous projection loss is calculated as the L2 distance between the predicted acoustic feature $\hat{\mathbf{a}}$ and the ground-truth acoustic feature $\mathbf{a}$, which is defined as:
@@
Discrete Projection: In general, the codec is designed with multi-layer quantizers to compress original speech into $L$-layer discrete indices $\mathbf{a}\in\mathcal{R}^{T\times L}$ at a low bitrate. Most LM-based work Wang et al. (2023b, c) stacks multiple LMs to predict discrete features, making the pipeline complicated and unsuitable for the streaming scenario. In contrast, StreamVoice adopts a streamlined multi-layer codec prediction method inspired by MQTTS Chen et al. (2023a). This method, free from temporal dependencies, can seamlessly integrate into the streaming process of the language model. To be specific, a single-layer transformer is used to model the heretical conditional distribution of codecs. As depicted in the right of Fig. 4, at time $t$, the transformer employs the ${}^{c}\mathbf{h}$ as the starting condition and sequentially generates $a^{l}_{t}$ from layer 1 to L. Remarkably, this generation process is independent of the preceding or the future $\mathbf{{}^{c}\mathbf{h}}$, rendering it well-suited for the demands of a streaming scenario. Notably, in the proposed StreamVoice, we mainly incorporate the discrete projection to achieve acoustic prediction. The discrete projection loss can be described as:
$\mathcal{L}_{Disc}=-\log{\prod^{T-1}_{t=0}p(a_{t}|\mathbf{a}_{1:t-1},^{m}\mathbf{s}_{1:t},t)}.$ (2)

### v5

@@ Abstract
Recent language model (LM) advancements have showcased impressive zero-shot voice conversion (VC) performance. However, existing LM-based VC models usually apply offline conversion from source semantics to acoustic features, demanding the complete source speech and limiting their deployment to real-time applications. In this paper, we introduce StreamVoice, a novel streaming LM-based model for zero-shot VC, facilitating real-time conversion given arbitrary speaker prompts and source speech. Specifically, to enable streaming capability, StreamVoice employs a fully causal context-aware LM with a temporal-independent acoustic predictor, while alternately processing semantic and acoustic features at each time step of autoregression which eliminates the dependence on complete source speech. To address the potential performance degradation from the incomplete context in streaming processing, we enhance the context-awareness of the LM through two strategies: 1) teacher-guided context foresight, using a teacher model to summarize the present and future semantic context during training to guide the model’s forecasting for missing context; 2) semantic masking strategy, promoting acoustic prediction from preceding corrupted semantic and acoustic input, enhancing context-learning ability. Notably, StreamVoice is the first LM-based streaming zero-shot VC model without any future look-ahead. Experiments demonstrate StreamVoice’s streaming conversion capability while achieving zero-shot performance comparable to non-streaming VC systems.
StreamVoice: Streamable Context-Aware Language Modeling for Real-time Zero-Shot Voice Conversion
Zhichao Wang1, Yuanzhe Chen2, Xinsheng Wang1, Lei Xie1111Corresponding author, Yuping Wang2 1Audio, Speech and Language Processing Group (ASLP@NPU) School of Computer Science, Northwestern Polytechnical University, Xi’an, China 2Douyin Vision Co., Ltd.
Figure 1: The concept of the streaming zero-shot VC employing the widely used recognition-synthesis framework Sun et al. (2016), where only the encoder of ASR is involved. StreamVoice is built on this popular paradigm.
@@ 1 Introduction
Voice conversion (VC) aims to transfer a speaker’s voice to that of another speaker without changing the linguistic content. This technique has been deployed in many real-world applications, such as movie dubbing, privacy protection, pronunciation correction, etc. With the help of neural semantic features, such as the bottleneck feature (BNF) from an automatic speech recognition (ASR) system, converting source speech from arbitrary speakers in the wild has been successfully achieved Sun et al. (2016). Meanwhile, converting to an arbitrary target speaker with only one utterance of this speaker, called zero-shot VC, has also been researched recently Qian et al. (2019); Wang et al. (2023c). However, most existing zero-shot VC models are designed for offline systems, which are insufficient to meet the recent growing demands of streaming capability in real-time VC applications, such as live broadcasting and real-time communication (RTC). In this study, we focus on the streaming zero-shot VC as illuminated in Fig. 1.
@@
Disentangling speech into different components, e.g., semantic content and speaker timbre, plays an important role in the zero-shot VC task Chou and Lee (2019); Wang et al. (2023d, 2021); Qian et al. (2019). Recently, benefiting from the powerful LM framework and the scaling up of training data, LM-based VC models Wang et al. (2023c); Yang et al. (2023); Zhu et al. (2023) with built-in in-context learning ability can learn the context relations between source and target speaker’s utterances to capture fine-grained speaker timbre, achieving impressive zero-shot VC performance. However, demanding the complete source speech utterance limits these LM-based VC models to real-time scenarios; thus, they can only be used in offline applications. While several non-LM-based methods Yang et al. (2022); Wang et al. (2023a) have been proposed for streaming zero-shot VC, the performance fails to generalize well to unseen speakers with high speaker similarity and speech naturalness, mainly due to the limited model capacity to scale up training data, and also the performance degradation caused by the missing future information in streaming scenario.
@@
Inspired by the success of LM-based models in zero-shot VC, we aim to explore the feasibility of LMs for the streaming VC scenario. An intuitive way is to follow the popular recognize-synthesis framework shown in Fig. 1, in which speech is represented in semantic BNF and acoustic features respectively extracted by a streaming ASR and an audio codec. Then, the LM-based VC model undertakes the transformation of semantic information into acoustic features with the target speaker’s timbre. However, the development of the LM-based model in streaming zero-shot VC is hampered by two primary challenges.
  1. $\bullet$
@@
    Streamable architecture: streaming models typically produce immediate output upon receiving current input without reliance on future time steps. Current LM-based VC models perform the conversion only when they get a full-utterance of source speech, which fails to meet the demands of streaming applications. The widely adopted multi-stage language modeling for multi-layer codec prediction introduces complexity to system design, posing a potential risk of cumulative errors. Additionally, the dependency models of the streaming pipeline also impact the design and performance of the VC model.
  2. $\bullet$
@@
    Performance gap: unlike non-streaming models, streaming models must process frame-wise or chunked input causally on the fly without future information, facing missing context and potential performance degradation. This missing hinders the streaming VC model from achieving high-quality conversion. In addition, as shown in Fig. 1, the VC model relies on the semantic feature BNF from ASR to achieve conversion, which makes semantic features very important. However, streaming ASR exhibits inferior performance compared to its non-streaming counterpart, leading to the BNF carrying low-quality semantic information but more speaker information. In addition to the inherent unavailable future reception, this low-quality semantic input makes achieving high-quality conversion more difficult. The goal of zero-shot VC amplifies the challenges faced by our streaming VC model.
@@
In this work, we propose StreamVoice, a streaming LM-based model for high-quality zero-shot VC. Specifically, StreamVoice has a streamable architecture that integrates a single-stage language model that casually generates acoustic codecs with the collaboration of an acoustic predictor. Alternating input of semantic and acoustic features at each time step ensures seamless streaming behavior. Two methods are introduced to enhance the context-awareness of the LM to mitigate the performance gap caused by missing contextual information. 1) We incorporate a teacher-guided context foresight, where the VC model is taught by a teacher non-streaming ASR to infer the present and future semantic information summarized by the teacher, which is then used to enhance the acoustic prediction. 2) To enhance the context learning from the input history, semantic masking encourages acoustic prediction from the preceding acoustic and corrupted semantic input, which also implicitly creates an information bottleneck to reduce the source speaker’s information.
@@
Experiments demonstrate StreamVoice’s ability to convert speech in a streaming manner with high speaker similarity for both seen and unseen speakers while maintaining performance comparable to non-streaming VC systems. As the first LM-based zero-shot VC model without any future look-ahead, the total pipeline only has 124 ms latency to perform the conversion, 2.4x faster than real-time on a single A100 GPU without engineering optimizations. Converted samples can be found in https://kerwinchao.github.io/StreamVoice/.
@@ 2 Related Works
Zero-shot Voice Conversion. Zero-shot VC imposes stringent demands on speech decoupling and capturing speaker timbre. Many studies specifically design many disentanglement approaches, incorporating intricate structures Chou and Lee (2019), loss functions Wang et al. (2021), and training strategies Ebbers et al. (2021), to achieve speech decoupling. Rather than embedding explicit disentanglement designs in VC training, some approaches Gu et al. (2021) leverage a speaker verification (SV) model for speaker representation, while linguistic content is extracted using ASR or self-supervised learning (SSL) models Sun et al. (2016); Choi et al. (2021). To enhance speaker timbre capturing, some fine-grained speaker modeling methods have also been explored Yin et al. (2021); Wang et al. (2023d). Recent successes of language models in generative tasks have prompted the exploration of LM-based models in zero-shot VC, yielding impressive results. Using the pre-trained model to decouple speech, the LM-based VC model Wang et al. (2023c); Yang et al. (2023); Zhu et al. (2023) captures fine-grained speaker timbre from the speaker prompt and then performs the conversion. However, current LM-based VC models are inapplicable to streaming scenarios, constraining their real-world utility. This paper addresses this gap by investigating the zero-shot capabilities of language models specifically tailored for streaming scenarios.
@@
Streaming Voice Conversion. Despite the high-quality conversion achieved by non-streaming VC models, their non-streamable structure and reliance on full-utterance input hamper them for real-time streaming applications. For streaming, causal processing and the structure of the streaming pipeline are crucial considerations. Streaming models are compelled to process frame-wise or chunked input on the fly, devoid of access to future information, leading to performance degradation compared to non-streaming counterparts. To address this, a common approach Hayashi et al. (2022); Kameoka et al. (2021); Ning et al. (2023, 2024) involves the integration of a teacher model to guide the training of the streaming model or the distillation of knowledge from a non-streaming model. Chen et al. Chen et al. (2023b) focus on selecting BNF with minimal semantic information loss through layer-wise analysis, while Chen et al. Chen et al. (2022) incorporate adversarial training to enhance the quality of semantic features. Beyond streaming VC, some efforts have recently been towards streaming zero-shot VC. For instance, VQMIVC Wang et al. (2021), designed for the non-streaming application, is modified to be streamable by Yang et al. Yang et al. (2022). ALO-VC Wang et al. (2023a) constructs a streaming system using an SV model, a streaming PPG extractor, and a pitch extractor. However, current streaming zero-shot VC, designed for low-resource devices, has limited model capacity with poor generalization to unseen speakers, leading to inferior similarity and naturalness. Motivated by LM’s successes in zero-shot VC, we design a streamable LM in streaming scenarios. To tackle distinctive challenges in streaming VC, we enhance the context awareness of the LM to improve conversion quality.
@@
Language Model-based Speech Generation. Recent advancements in LMs within natural language processing have showcased potent generation capabilities, influencing the development of LMs in speech generation. By employing codec Zeghidour et al. (2021) or other SSL models Chung et al. (2021), speech and audio can be efficiently tokenized into discrete units, facilitating low-bitrate audio representation and semantic extraction. This progress allows speech generation to utilize LM frameworks seamlessly. Taking audio generation as a conditional language modeling task, AudioLM Borsos et al. (2023) employs hierarchical language modeling for acoustic prediction from coarse to fine units. VALL-E Wang et al. (2023b) and SpearTTS Kharitonov et al. (2023) extend LMs for zero shot-TTS, which can clone a human’s voice with prompt tokens from a short recording. For zero-shot VC, LM-VC Wang et al. (2023c) employs task-oriented optimizations to this task. And some studies Zhu et al. (2023); Yang et al. (2023) leverage multitask objectives and datasets, achieving high-quality conversion. Despite this progress, existing LM-based VC models usually apply offline processing, demanding complete utterance from the source speech, which hinders their suitability for real-time streaming applications. In contrast to prior studies, we explore the zero-shot capability of the LM-based VC for streaming scenarios. With the enhancement of context awareness, the proposed LM-based VC model achieves results comparable to non-streaming LM-based VC.
@@ 3.1 Overview
As shown in Fig. 2, the development of StreamVoice follows the recognition-synthesis framework. In this framework, speech is first represented as semantic features $\mathbf{s}=\{s_{1},s_{2},...s_{T_{s}}\}$ and acoustic features $\mathbf{a}=\{a_{1},a_{2},...,a_{T_{a}}\}$ by a pre-trained streaming ASR model and a speech codec model respectively. Here, $T_{s}$ and $T_{a}$ denote the sequence length. Before inputting to StreamVoice, $\mathbf{s}$ and $\mathbf{a}$ are aligned to the same length $T$. StreamVoice incorporates a context-aware language model and an acoustic predictor to perform a single language modeling process. With the semantic and acoustic features $\{\mathbf{\tilde{s}},\mathbf{\tilde{a}}\}$ of speech from the target speaker as speaker prompt, the LM leverages the semantic information $\mathbf{s}_{1:t}$ of source speech to autoregressively predict the hidden output ${}^{c}\mathbf{h}$. In each autoregression time-step of the LM, the acoustic predictor transforms the hidden output ${}^{c}\mathbf{h}$ to the codec feature $\mathbf{\hat{a}}$ of the converted speech. Finally, the codec model reconstructs the waveform from the predicted codec feature. In the following sections, we will introduce how to build a streamable LM for VC and how to ensure the high-quality conversation of this streaming VC.
Figure 3: The architecture for context-aware LM.
@@ 3.2 Streamable Architecture
To perform streaming VC, a streamable architecture is necessary. In StreamVoice, the language model is carefully designed to perform full causal processing in the VC task, and the acoustic predictor is designed to achieve frame-wise prediction without dependency on temporal information.
@@ 3.2.1 Fully Casual Language Model
To be specific, the semantic feature $\mathbf{s}$ obtained via an ASR model comprises a sequence of embeddings, denoted as $\{s_{1},s_{2},...,s_{T}\}$. On the other hand, the codec tokens obtained from an $L$-layer codec are discrete units represented by $\mathbf{a}\in\mathcal{R}^{T\times L}$. To obtain the acoustic embedding sequence, the codec tokens from each layer undergo separate embedding into the embedding space, and then they are concatenated along the embedding dimension, resulting in the fused acoustic embedding. Both the fused acoustic embedding and semantic features are transformed into the same dimension using linear layers. Subsequently, they are alternately inputted into the language model, forming the cross-embedding.
Figure 4: The architecture for acoustic predictor. Our system can support continuous or discrete projection.
@@ 3.2.2 Acoustic Predictor
As the preceding LM has essentially encoded content and speaker into its output ${}^{c}\mathbf{h}$, the acoustic predictor can be designed in temporal irrelevant to transform ${}^{c}\mathbf{h}$ into acoustic codec space, which means the predictor can be easily applied in the streaming scenario. Given that the speech can be represented in acoustic features by neural codec in either continuous or discrete forms, we investigate the incorporation of both features in StreamVoice, which are performed by continuous projection and discrete projection, respectively.
Continuous Projection. Following Shen et al. Shen et al. (2024), the $D$-dimensional quantized latent vector $\mathbf{a}\in\mathcal{R}^{T\times D}$ encoded by the codec model is used as the continuous acoustic representation. The prediction of the continuous representation involves employing a stack of linear layers, as shown in Fig. 4. The continuous projection loss is calculated as the L2 distance between the predicted acoustic feature $\hat{\mathbf{a}}$ and the ground-truth acoustic feature $\mathbf{a}$, which is defined as:
@@
Discrete Projection. In general, the codec is designed with multi-layer quantizers to compress original speech into $L$-layer discrete indices $\mathbf{a}\in\mathcal{R}^{T\times L}$ at a low bitrate. Most LM-based work Wang et al. (2023b, c) stacks multiple LMs to predict discrete features, making the pipeline complicated and unsuitable for the streaming scenario. In contrast, StreamVoice adopts a streamlined multi-layer codec prediction method inspired by MQTTS Chen et al. (2023a). This method, free from temporal dependencies, can seamlessly integrate into the streaming process of the language model. Specifically, a single-layer transformer is used to model the hierarchical conditional distribution of codecs. As depicted in the right of Fig. 4, at time $t$, the transformer employs the ${}^{c}\mathbf{h}$ as the starting condition and sequentially generates $a^{l}_{t}$ from layer 1 to L. Remarkably, this generation process is independent of the preceding or the future $\mathbf{{}^{c}\mathbf{h}}$, rendering it well-suited for the demands of a streaming scenario. Notably, in the proposed StreamVoice, we mainly incorporate the discrete projection to achieve acoustic prediction. The discrete projection loss can be described as:
$\mathcal{L}_{Disc}=-log{\prod_{t=1}^{T}\prod_{l=1}^{L}p(a^{l}_{t}|\mathbf{a}_{1:t-1},^{m}\mathbf{s}_{1:t},t,a^{1:l-1}_{t})}.$ (2)